OpenStudy (anonymous):

6 ( n - 3 ) = 5 ( n + 7 ) solving proportions

OpenStudy (anonymous):

n=53

OpenStudy (anonymous):



OpenStudy (anonymous):

\[6(n-3)=5(n+7)\] \[6n-18=5n+35\] \[6n-5n=35+18\] \[n=53\]

OpenStudy (anonymous):

6n-18=5n+35 6n-5n=35+18 n=53

OpenStudy (anonymous):

thanks